OpenStudy (anonymous):

Inverse laplace ((s+1)/(s^2))(e^-s)

OpenStudy (anonymous):

Please do help

OpenStudy (tkhunny):

Have you considered \(\dfrac{s+1}{s^{2}} = \dfrac{1}{s}+\dfrac{1}{s^{2}}\)

OpenStudy (anonymous):

not yet is that so?

OpenStudy (tkhunny):

I think addition of fractions still works, doesn't it?

OpenStudy (anonymous):

its a inverse laplace transform

OpenStudy (tkhunny):

Yes. It's okay to use algebra to simplify your task.

OpenStudy (tkhunny):

\(\mathscr{L}^{-1}\left(\dfrac{s+1}{s^{2}}\cdot e^{-s}\right) = \mathscr{L}^{-1}\left(\dfrac{e^{-s}}{s}\right) + \mathscr{L}^{-1}\left(\dfrac{e^{-s}}{s^{2}}\right)\) and one or both of those may look more familiar.

OpenStudy (anonymous):

is that the final answer?

OpenStudy (tkhunny):

No. There is no inverse, yet. Just algebra. You don't recognize those and being sufficiently common to include in a table?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

how about \[\frac{6 (s+1) }{ s^4 }\]

OpenStudy (anonymous):

its also an inverse laplace

OpenStudy (tkhunny):

How did you get "s"? Normally, it moves to 't'. It is very important that you know \(\mathscr{L}^{-1}\left(\dfrac{e^{-s}}{s}\right) = \Phi(t-1)\)

OpenStudy (tkhunny):

t wouldn't hurt also to know that \(\mathscr{L}^{-1}\left(\dfrac{e^{-s}}{s^{2}}\right) = (t-1)\cdot \Phi(t-1)\) Have you seen those things?

OpenStudy (anonymous):

nope

OpenStudy (tkhunny):

Okay, I'm a little stunned by that. It's usually Page 1 in Laplace Transform world. Please look up the Heaviside Step Function and its Laplace Transform. Really, why would you be given this material if you are not also introduced to the tools to achieve or to understand the solution? Something seriously wrong.

OpenStudy (anonymous):

yes really I really dont understand this thing thats whay I beg for help thanks for helping.

OpenStudy (tkhunny):

Well, we won't get far without Heaviside. Let's try the easy one. \(\mathscr{L}^{-1}\left(\dfrac{s+1}{s^{4}}\right) = \mathscr{L}^{-1}\left(\dfrac{1}{s^{3}}\right) + \mathscr{L}^{-1}\left(\dfrac{1}{s^{4}}\right)\) Now what?

OpenStudy (anonymous):

54

OpenStudy (tkhunny):

No. Not even close. Why are you in this material if you have not actually been introduced to it? Please go have a very sincere chat with your academic advisor. You will spend a very frustrated academic year if you continue on this path.